payton:

3x - 7y = -14 x=2y-3 Which equation is substituted correctly? 2y - 3 = 0 3x - 7(2y-3) = -14 3(2y-3) = -14 3(2y-3) - 7y = -14

TheSmartOne:

Feel like there's something missing from this question

ramen:

Same

MrMudd183:

that's what i was thinking

payton:

x=2y-3

payton:

i forgot that part at the top

TheSmartOne:

Does it give you any other equation?

payton:

3x - 7y = -14

TheSmartOne:

Yeeeaaaaahhhh Should have probably included that too xD So we have a x= 2y-3 And we have that equation So you can substitute one into the other Here's an example so you can see how a = 3b 2a + 2b = 4 Because our first equation is a = 3b That means since they're equal that they're interchangable so you go from 2a + 2b = 4 To 2(3b) + 2b = 4

TheSmartOne:

Does that make sense?

payton:

yea sortta

TheSmartOne:

So were you able to figure it out?